OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)?

OpenStudy (anonymous):

y=2x+5

OpenStudy (anonymous):

(y-1) = 2* (x+2)

OpenStudy (anonymous):

wrong

OpenStudy (amistre64):

y-1=\(\cfrac{13-1}{4--2}\)(x--2)